# Ideals, Gröbner Bases, and PCPs

### v1

@@ 2.1 PCPs
Let $\mathbb{F}_{q}$ be a field of characteristic $\neq 2$ and let $V\subseteq\mathbb{F}_{q}^{m}$ have extension degree $d$ and Gröbner complexity $k$. Then 33-$\mathsf{COLOR}$ on vertex set $V$ is in $\mathsf{PCP}_{1,1/2}[c(k+m)\log q,\;\ell,\;cd\log q]$ with proofs of size $q^{c(k+m)}$.
@@
In Theorem 2.3, the assumption on characteristic $\neq 2$ is mostly for clarity in the proofs. We assign the vertices colors from the set $\left\{-1,0,1\right\}$, and these are three distinct colors only if the field is of characteristic $\neq 2$. For fields of characteristic $2$ and with more than $3$ elements, one could use colors $\left\{a,b,c\right\}$, where $a,b,$ and $c$ are three distinct elements from the field. The proof is essentially the same.
@@
In Corollary 4.4 we show that if $V=H^{m}$ for some $H\subseteq\mathbb{F}_{q}$ then it has extension degree $(|H|-1)\cdot m$ and Gröbner complexity $m$. Taking $q\geq c\log^{6}n$ a power of $3$, $V=H^{m}$ for some subset $H\subseteq\mathbb{F}_{q}$ of size $\log n$ and $m=\log n/\log\log n$, we get the desired bounds. ∎
@@
In Example 4.2.2 we show that if $V=\{0,1\}^{m}_{\leq 1}:=\left\{(a_{1},\ldots,a_{m})\in\{0,1\}^{m}|\sum_{i=1}^{m}a_{i}\leq 1\right\}$ is the set of Boolean points in $\mathbb{F}_{q}^{m}$ of Hamming weight at most $1$, then $V$ has extension degree $1$ and Gröbner complexity $\mathcal{O}(m^{2})$. Picking $q$ to be a large constant, and $V=\{0,1\}^{m}_{\leq 1}$ we get the desired bounds. ∎
Note that the above roughly matches the parameters of the Hadamard PCP of [ALMSS] with a completely different proof!
@@
In Corollary 4.5 we show that if $V=\left(\{0,1\}^{m}_{\leq 1}\right)^{c}$, then $V$ has extension degree $c$ and Gröbner complexity $\mathcal{O}(cm^{2})$. Given $\varepsilon>0$ picking $c=\mathcal{O}(\frac{1}{\varepsilon})$, $q=\operatorname{poly}(1/\varepsilon)$, $m=n^{\mathcal{O}(\varepsilon)}$, and $V=\left(\{0,1\}^{m}_{\leq 1}\right)^{c}$ we get the desired bounds. ∎
@@ 3 Preliminaries
Theorem 3.1 (Polynomial Distance Lemma).
[ore1922hohere, DL78, Schwartz80, Zippel79]. Fix a field $\mathbb{F}_{q}$. For every degree parameter $d\in\mathbb{N}$ with $d\leq q$ and for every non-zero polynomial $P\in\mathbb{F}_{q}[x_{1},\ldots,x_{m}]$, we have:
@@
An immediate and useful corollary of Theorem 3.1 is the following: If two degree $\leq d$ polynomials $P$ and $Q$ agree on strictly more than $d/q$-fraction of $\mathbb{F}_{q}^{m}$, then $P=Q$.
@@ 3.1 Low-degree Testing
In this subsection, we discuss the standard point-vs-line test for low-degree testing from [ALMSS]. We start by recalling the test and state its properties in Theorem 3.5. In the following discussion, we will switch between a function $f:\mathbb{F}_{q}^{m}\to\mathbb{F}_{q}$ and its evaluation vector $f\in(\mathbb{F}_{q})^{q^{m}}$, as both are equivalent.
1 Input: Degree parameter $d$, string $\mathbf{a},\mathbf{b}\in\mathbb{F}_{q}^{m}$, element $t\in\mathbb{F}_{q}^{\times}$, and oracle access to $(f,f^{\prime})$ where $f\in(\mathbb{F}_{q})^{q^{m}}$ and $f^{\prime}\in(\mathbb{F}_{q}^{d+1})^{q^{2m}}$ 2 Query $f^{\prime}[({\bf a},{\bf b})]$ and query $f[{{\bf a}+t{\bf b}}]$ // Two queries to $(f,f^{\prime})$ 3 4 if $f^{\prime}[({\bf a},{\bf b})](t)\neq f[{{\bf a}+t{\bf b}}]$ // Running time is $\operatorname{poly}(m,d)$ 5 then 6  7  3mm return Reject else 8    return Accept 9  10 Algorithm 1 Low-Degree Test $\mathcal{LDT}^{(\cdot)}_{\cdot}$
Theorem 3.5 (Low-degree Testing (see for instance [ALMSS, Theorem A5])).
@@
For low-degree testing, there has been a long line of work on achieving better parameters in terms of field size and soundness guarantee. We refer the interested reader to [HKSS-LDT, Section 1] for a detailed overview of the results of low-degree testing and also for the state-of-the-art parameters (see [HKSS-LDT, Theorem 1.2]). We use the low-degree testing from [ALMSS] because the algorithm and analysis are done using the lines table.
@@ 3.2 Local Correction of Low-Degree Polynomials
In this subsection, we discuss the local correction/self-correction algorithm for degree $d$ polynomials over $\mathbb{F}_{q}^{m}$ from [ALMSS]. We first describe the local corrector and then analyze it in Theorem 3.7. In the following discussion, we will switch between a function $f:\mathbb{F}_{q}^{m}\to\mathbb{F}_{q}$ and its evaluation vector $f\in(\mathbb{F}_{q})^{q^{m}}$, as both are equivalent.
1 Input: Degree parameter $d$, evaluation point $\mathbf{a}\in\mathbb{F}_{q}^{m}$, string $\mathbf{b}\in\mathbb{F}_{q}^{m}$, element $t\in\mathbb{F}_{q}^{\times}$, and oracle access to $(f,f^{\prime})$ where $f\in(\mathbb{F}_{q})^{q^{m}}$ and $f^{\prime}\in(\mathbb{F}_{q}^{d+1})^{q^{2m}}$ 2 3 Query $f^{\prime}[(\mathbf{a},\mathbf{b})](t)$ and query $f[{{\bf a}+t{\bf b}}]$ // Two queries to $(f,f^{\prime})$ 4 5 if $f^{\prime}[({\mathbf{a},\mathbf{b}})](t)\neq f[\ell_{\mathbf{a},\mathbf{b}}(t)]$ // Running time is $\operatorname{poly}(m,d)$ 6 then 7  8  9 return Reject return $f^{\prime}[(\mathbf{a},\mathbf{b})](0)$ Algorithm 2 Local Corrector $\mathcal{LC}^{(\cdot)}_{\cdot}$
Theorem 3.7 (Local Correction (see e.g. [ALMSS], Proposition 7.2.2.1)).
@@ 4 Gröbner Generating Sets
Definition 4.2 (Gröbner Complexity).
For an ideal $\mathbb{I}\subseteq\mathbb{F}_{q}[x_{1},\ldots,x_{m}]$ we say that a finite set $\mathfrak{G}\subseteq\mathbb{I}$ is a Gröbner generating set of $\mathbb{I}$, if every polynomial $P\in\mathbb{I}$ can be written as follows:
@@
For a set $V\subseteq\mathbb{F}_{q}^{m}$, let $\mathbb{I}(V)$ denote the ideal of polynomials that vanish on $V$. We define the Gröbner complexity of $V$ to be the cardinality of the smallest Gröbner generating set of $\mathbb{I}(V)$.
The naming convention comes from the fact that a Gröbner basis in a graded ordering is a Gröbner generating set, as we show in Appendix A.
@@
Furthermore, for any function $f:H\to\mathbb{F}_{q}$ we can find a degree |H|1|H|-1 polynomial extending $f$. It follows $H$ has Gröbner complexity 1, and extension degree |H|1|H|-1.
@@
$\left(x_{1}^{2}-x_{1},\dots,x_{n}^{2}-x_{n},x_{1}x_{2},\dots,x_{n-1}x_{n}\right)$
is a Gröbner generating set. To see this, note that any polynomial $P$ can be written on the form
$P(\mathbf{x})=\sum_{i}^{n}\;h_{i}(\mathbf{x})\cdot(x_{i}^{2}-x_{i})\;+\;\sum_{i,j}^{n}\;g_{i,j}(\mathbf{x})\cdot(x_{i}x_{j})\;+\;\ell(\mathbf{x})$
@@
It follows that $\{0,1\}^{n}_{\leq 1}$ has Gröbner complexity at most $\frac{n(n+1)}{2}$ and extension degree $1$.
The following lemma shows that we can upper-bound both the Gröbner complexity and the extension degree of Cartesian products.
Lemma 4.3 (Subadditivity of Gröbner complexity and extension degree).
@@
    if $\mathfrak{G}_{1},\mathfrak{G}_{2}$ are Gröbner generating sets for $\mathbb{I}(V_{1}),\mathbb{I}(V_{2})$ respectively, then $\mathfrak{G}_{1}\cup\mathfrak{G}_{2}$ is a Gröbner generating set for $\mathbb{I}(V_{1}\times V_{2})$.
@@
In particular, both the Gröbner complexity and extension degree are subadditive under Cartesian products.
@@
is a polynomial of degree at most $\deg(P)$ vanishing on $V$. It follows that a subset $\mathfrak{G}\subseteq\mathbb{F}_{q}[\mathbf{x}]$ is a Gröbner generating set of $\mathbb{I}(V)$ if and only if every polynomial $P$ can be written in the form
@@
where each summand has degree at most $\deg(P)$.
@@
Then $S_{12}$ is a basis for functions $V_{1}\times V_{2}\to\mathbb{F}_{q}$ with the above mentioned property. We will show that $\mathfrak{G}_{1}\cup\mathfrak{G}_{2}$ is a Gröbner generating set for $V_{1}\times V_{2}$. Let $m_{1}(\mathbf{x})m_{2}(\mathbf{y})$ be a monomial in $\mathbb{F}_{q}[\mathbf{x},\mathbf{y}]$, We can apply the above property to each monomial separately to get
@@
We see this gives a representation of the monomial as a linear combination of terms from $S_{12}$ and $\mathfrak{G}_{1}\cup\mathfrak{G}_{2}$ of degree at most $\deg(m_{1})+\deg(m_{2})$. Since we can write a polynomial as a sum of monomials, we get that $\mathfrak{G}_{1}\cup\mathfrak{G}_{2}$ is a Gröbner generating set. This proves the first item of the lemma.
@@
For any subset $H\subseteq\mathbb{F}_{q}$, $H^{m}\subseteq\mathbb{F}_{q}^{m}$ has Gröbner complexity at most $m$ and extension degree at most $\left(|H|-1\right)\cdot m$.
@@
The subset $\left(\{0,1\}^{n/c}_{\leq 1}\right)^{c}\in\mathbb{F}_{q}^{n}$ has Gröbner complexity at most $\frac{(n^{2}+nc)}{2c}$ and extension degree at most $c$.
@@
The following two lemmas show that given a subset $V$ of $\mathbb{F}_{q}^{m}$, we can efficiently compute its extension degree, find a polynomial extending a function $f:V\to\mathbb{F}_{q}$, and compute a minimal Gröbner basis.
@@
For every $q,m\in\mathbb{N}$, there exists:
@@
Lemma 4.7 (Computing a smallest Gröbner generating set).
For every $q,m\in\mathbb{N}$, there exists an algorithm which takes a set of $n$ points $V$ in $\mathbb{F}_{q}^{m}$ as input and gives a minimal (w.r.t. size) Gröbner generating set of $\mathbb{I}(V)$ as output in the monomial basis in time $\operatorname{poly}\left(n,\log q,\binom{m+d+1}{m}\right)$, where $d$ is the extension degree of $V$.
@@
We will then construct the subsets of the minimal Gröbner generating set $\mathfrak{G}_{i}$ inductively. We first define $\mathfrak{G}_{0}=\emptyset$, and then set $\mathfrak{G}_{i}$ to be any basis of the quotient space $\mathbb{I}(V)_{\leq i}/L_{i}$, where
@@
We will now show any set $\mathfrak{G}$ is a minimal Gröbner generating set if and only if $\mathfrak{G}_{i}$ is a basis for the quotient space $\mathbb{I}(V)_{\leq i}/L_{i}$. To see this, if every $\mathfrak{G}_{i}$ is a spanning set for $\mathbb{I}(V)_{\leq i}/L_{i}$, then any polynomial $P\in\mathbb{I}(V)_{\leq i}$ can be written as
@@
where $c_{g}$ are constants and $\deg(h_{g}\cdot g)\leq i$. This is the condition for $\mathfrak{G}$ being a Gröbner generating set, so $\mathfrak{G}$ is a Gröbner generating set if and only if each $\mathfrak{G}_{i}$ spans $\mathbb{I}(V)_{\leq i}/L_{i}$.
@@
From the above, we see that a minimal Gröbner generating set does not contain any degree $i$ polynomials if

### v2

@@ 2.1 PCPs
Let $\mathbb{F}_{q}$ be a field of characteristic $\neq 2$ and let $V\subseteq\mathbb{F}_{q}^{m}$ have extension degree $d$ and Macaulay complexity $k$. Then 33-$\mathsf{COLOR}$ on vertex set $V$ is in $\mathsf{PCP}_{1,1/2}[c(k+m)\log q,\;\ell,\;cd\log q]$ with proofs of size $q^{c(k+m)}$.
@@
In Theorem 2.3, the assumption on characteristic $\neq 2$ is mostly for clarity in the proofs. We assign the vertices colors from the set $\left\{-1,0,1\right\}$, and these are three distinct colors only if the field is of characteristic $\neq 2$. For fields of characteristic $2$ and with at least $3$ elements, one could use colors $\left\{a,b,c\right\}$, where $a,b,$ and $c$ are three distinct elements from the field. The proof is essentially the same.
With this setup, we first show the following two results.
@@
In Corollary 4.5 we show that if $V=H^{m}$ for some $H\subseteq\mathbb{F}_{q}$ then it has extension degree $(|H|-1)\cdot m$ and Macaulay complexity $m$. Taking $q\geq c\log^{6}n$ a power of $3$, $V=H^{m}$ for some subset $H\subseteq\mathbb{F}_{q}$ of size $\log n$ and $m=\log n/\log\log n$, we get the desired bounds. ∎
@@
In Example 4.3.2 we show that if $V=\{0,1\}^{m}_{\leq 1}:=\left\{(a_{1},\ldots,a_{m})\in\{0,1\}^{m}|\sum_{i=1}^{m}a_{i}\leq 1\right\}$ is the set of Boolean points in $\mathbb{F}_{q}^{m}$ of Hamming weight at most $1$, then $V$ has extension degree $1$ and Macaulay complexity $\mathcal{O}(m^{2})$. Picking $q$ to be a large constant, and $V=\{0,1\}^{m}_{\leq 1}$ we get the desired bounds. ∎
Note that the above roughly matches the parameters of the Hadamard PCP of [2] with a completely different proof!
@@
In Corollary 4.6 we show that if $V=\left(\{0,1\}^{m}_{\leq 1}\right)^{c}$, then $V$ has extension degree $c$ and Macaulay complexity $\mathcal{O}(cm^{2})$. Given $\varepsilon>0$ picking $c=\mathcal{O}(\frac{1}{\varepsilon})$, $q=\operatorname{poly}(1/\varepsilon)$, $m=n^{\mathcal{O}(\varepsilon)}$, and $V=\left(\{0,1\}^{m}_{\leq 1}\right)^{c}$ we get the desired bounds. ∎
@@ 3 Preliminaries
Fact 3.1 (Polynomial Distance Lemma).
[18, 8, 20, 22]. Fix a field $\mathbb{F}_{q}$. For every degree parameter $d\in\mathbb{N}$ with $d\leq q$ and for every non-zero polynomial $P\in\mathbb{F}_{q}[x_{1},\ldots,x_{m}]$, we have:
@@
An immediate and useful corollary of 3.1 is the following: If two degree $\leq d$ polynomials $P$ and $Q$ agree on strictly more than $d/q$-fraction of $\mathbb{F}_{q}^{m}$, then $P=Q$.
@@ 3.1 Low-degree Testing
In this subsection, we discuss the standard point-vs-line test for low-degree testing from [2]. We start by recalling the test and state its properties in Theorem 3.5. In the following discussion, we will switch between a function $f:\mathbb{F}_{q}^{m}\to\mathbb{F}_{q}$ and its evaluation vector $f\in(\mathbb{F}_{q})^{q^{m}}$, as both are equivalent.
1 Input: Degree parameter $d$, randomness $\mathbf{a},\mathbf{b}\in\mathbb{F}_{q}^{m}$, $t\in\mathbb{F}_{q}^{\times}$, and oracle access to $(f,f^{\prime})$ where $f\in(\mathbb{F}_{q})^{q^{m}}$ and $f^{\prime}\in(\mathbb{F}_{q}^{d+1})^{q^{2m}}$ 2 Query $f^{\prime}[({\bf a},{\bf b})]$ and query $f[{{\bf a}+t{\bf b}}]$ // Two queries to $(f,f^{\prime})$ 3 4 if $f^{\prime}[({\bf a},{\bf b})](t)\neq f[{{\bf a}+t{\bf b}}]$ // Running time is $\operatorname{poly}(m,d)$ 5 then 6  7  3mm return Reject else 8    return Accept 9  10 Algorithm 1 Low-Degree Test $\mathcal{LDT}^{(\cdot)}_{\cdot}$
Theorem 3.5 (Low-degree Testing (see for instance [2, Theorem A5])).
@@
For low-degree testing, there has been a long line of work on achieving better parameters in terms of field size and soundness guarantee. We refer the interested reader to [13, Section 1] for a detailed overview of the results of low-degree testing and also for the state-of-the-art parameters (see [13, Theorem 1.2]). We use the low-degree testing from [2] because the algorithm and analysis are done using the lines table.
@@ 3.2 Local Correction of Low-Degree Polynomials
In this subsection, we discuss the local correction/self-correction algorithm for degree $d$ polynomials over $\mathbb{F}_{q}^{m}$ from [2]. We first describe the local corrector and then analyze it in Theorem 3.7. In the following discussion, we will switch between a function $f:\mathbb{F}_{q}^{m}\to\mathbb{F}_{q}$ and its evaluation vector $f\in(\mathbb{F}_{q})^{q^{m}}$, as both are equivalent.
1 Input: Degree parameter $d$, evaluation point $\mathbf{a}\in\mathbb{F}_{q}^{m}$, randomness $\mathbf{b}\in\mathbb{F}_{q}^{m}$, $t\in\mathbb{F}_{q}^{\times}$, and oracle access to $(f,f^{\prime})$ where $f\in(\mathbb{F}_{q})^{q^{m}}$ and $f^{\prime}\in(\mathbb{F}_{q}^{d+1})^{q^{2m}}$ 2 3 Query $f^{\prime}[(\mathbf{a},\mathbf{b})](t)$ and query $f[{{\bf a}+t{\bf b}}]$ // Two queries to $(f,f^{\prime})$ 4 5 if $f^{\prime}[({\mathbf{a},\mathbf{b}})](t)\neq f[\ell_{\mathbf{a},\mathbf{b}}(t)]$ // Running time is $\operatorname{poly}(m,d)$ 6 then 7  8  9 return Reject return $f^{\prime}[(\mathbf{a},\mathbf{b})](0)$ Algorithm 2 Local Corrector $\mathcal{LC}^{(\cdot)}_{\cdot}$
Theorem 3.7 (Local Correction (see e.g. [2], Proposition 7.2.2.1)).
@@ 4 Macaulay Bases
Definition 4.2 (Macaulay Complexity).
For an ideal $\mathbb{I}\subseteq\mathbb{F}_{q}[x_{1},\ldots,x_{m}]$ we say that a finite set $\mathfrak{G}\subseteq\mathbb{I}$ is a Macaulay basis of $\mathbb{I}$, if every polynomial $P\in\mathbb{I}$ can be written as follows:
@@
For a set $V\subseteq\mathbb{F}_{q}^{m}$, let $\mathbb{I}(V)$ denote the ideal of polynomials that vanish on $V$. We define the Macaulay complexity of $V$ to be the cardinality of the smallest Macaulay basis of $\mathbb{I}(V)$.
A few remarks about Macaulay bases follow below.
Remark 4.3.
Note that a Macaulay basis need not be a “minimal” one, i.e., it may be possible to add polynomials to a Macaulay basis while still keeping it a Macaulay basis. We also abuse notation slightly and refer to a Macaulay basis of a variety $V$ as a Macaulay basis of the corresponding ideal $\mathbb{I}(V)$. A Gröbner basis in a graded ordering is always a Macaulay basis. However a minimal Gröbner basis is not necessarily a minimal Macaulay basis, see Remark A.4.
@@
Furthermore, for any function $f:H\to\mathbb{F}_{q}$ we can find a degree |H|1|H|-1 polynomial extending $f$. It follows $H$ has Macaulay complexity 1, and extension degree |H|1|H|-1.
@@
$\left\{x_{i}^{2}-x_{i}\right\}_{1\leq i\leq n}\cup\left\{x_{1}x_{2},\dots,x_{n-1}x_{n}\right\}_{1\leq i<j\leq n}$
is a Macaulay basis. To see this, note that any polynomial $P$ can be written in the form
$P(\mathbf{x})=\sum_{i}^{n}\;h_{i}(\mathbf{x})\cdot(x_{i}^{2}-x_{i})\;+\;\sum_{i<j}^{n}\;g_{i,j}(\mathbf{x})\cdot(x_{i}x_{j})\;+\;\ell(\mathbf{x})$
@@
It follows that $\{0,1\}^{n}_{\leq 1}$ has Macaulay complexity at most $\frac{n(n+1)}{2}$ and extension degree $1$.
The above example can be extended to Hamming balls of radius greater than 1.
Example 4.3.3.
Let $\{0,1\}^{m}_{\leq c}\subseteq\mathbb{F}_{q}^{m}$ be the subset of $\{0,1\}^{m}$ consisting of points with Hamming weight at most $c$.
$\left\{x_{i}^{2}-x_{i}\right\}_{1\leq i\leq n}\cup\left\{x_{i_{1}}x_{i_{2}}\cdots x_{i_{c+1}}\right\}_{1\leq i_{1}<\cdots<i_{c+1}\leq n}$
is a Macaulay basis. To see this, any polynomial can be written on the form
$P(\mathbf{x})=\sum_{i}^{n}\;h_{i}(\mathbf{x})\cdot(x_{i}^{2}-x_{i})\;+\;\sum_{i_{1}<\ldots<i_{c+1}}^{n}\;g_{i,j}(\mathbf{x})\cdot(x_{i_{1}}x_{i_{2}}\cdots x_{i_{c+1}})\;+\;m(\mathbf{x})$
where $m(\mathbf{x})$ is a multilinear polynomial of degree at most $c$, and $\deg(h_{i})\leq\deg(P)-2,\deg(g_{i,j})\leq\deg(P)+1-c$.
Since the first two summands vanish on $\{0,1\}^{m}_{\leq c}$, any polynomial is therefore equivalent to a multilinear polynomial of degree at most $c$ on $P$, so multilinear polynomials span the set of functions on $\left\{0,1\right\}_{\leq c}^{m}$. However, since both the size of $\{0,1\}^{m}_{\leq c}$ and the number of multilinear monomials of degree at most $c$ is $\sum_{i=0}^{c}\binom{m}{i}$, it follows that a multilinear polynomial can only vanish on $V$ if it is 0. It follows that $f_{|V}=0$ if and only if $m(\mathbf{x})=0$ if and only if $f$ lies in the Macaulay basis.
This set then has Macaulay complexity $m+{m\choose c+1}=\Theta_{c}(m^{c+1})$ and extension degree $c$.
The following lemma shows that we can upper-bound both the Macaulay complexity and the extension degree of Cartesian products.
Lemma 4.4 (Subadditivity of Macaulay complexity and extension degree).
@@
    if $\mathfrak{G}_{1},\mathfrak{G}_{2}$ are Macaulay bases for $\mathbb{I}(V_{1}),\mathbb{I}(V_{2})$ respectively, then $\mathfrak{G}_{1}\cup\mathfrak{G}_{2}$ is a Macaulay basis for $\mathbb{I}(V_{1}\times V_{2})$.
@@
In particular, both the Macaulay complexity and extension degree are subadditive under Cartesian products.
@@
is a polynomial of degree at most $\deg(P)$ vanishing on $V$. It follows that a subset $\mathfrak{G}\subseteq\mathbb{F}_{q}[\mathbf{x}]$ is a Macaulay basis of $\mathbb{I}(V)$ if and only if every polynomial $P$ can be written in the form
@@
where each non-zero summand has degree at most $\deg(P)$.
@@
Then $S_{12}$ is a basis for functions $V_{1}\times V_{2}\to\mathbb{F}_{q}$ with the above mentioned property. We will show that $\mathfrak{G}_{1}\cup\mathfrak{G}_{2}$ is a Macaulay basis for $V_{1}\times V_{2}$. Let $m_{1}(\mathbf{x})m_{2}(\mathbf{y})$ be a monomial in $\mathbb{F}_{q}[\mathbf{x},\mathbf{y}]$, We can apply the above property to each monomial separately to get
@@
We see this gives a representation of the monomial as a linear combination of terms from $S_{12}$ and $\mathfrak{G}_{1}\cup\mathfrak{G}_{2}$ of degree at most $\deg(m_{1})+\deg(m_{2})$. Since we can write a polynomial as a sum of monomials, we get that $\mathfrak{G}_{1}\cup\mathfrak{G}_{2}$ is a Macaulay basis. This proves the first item of the lemma.
@@
For any subset $H\subseteq\mathbb{F}_{q}$, $H^{m}\subseteq\mathbb{F}_{q}^{m}$ has Macaulay complexity at most $m$ and extension degree at most $\left(|H|-1\right)\cdot m$.
@@
For $c\in N$ that divides $m$, the subset $\left(\{0,1\}^{m/c}_{\leq 1}\right)^{c}\subseteq\mathbb{F}_{q}^{m}$ has Macaulay complexity at most $\frac{m^{2}+cm}{2c}$ and extension degree at most $c$.
@@
Remark 4.7.
The sets given by Corollary 4.6 and Example 4.3.3 have similar extension degree (i.e., $c$) and size (i.e., $\Theta_{c}(m^{c})$). We note however that the Macaulay complexity bound for the latter family grows with $c$ as $\Theta(m^{c+1})$ (unlike the former family which is always $O(m^{2})$), and so the set given by Example 4.3.3 does not yield PCPs matching the parameters we get by using Corollary 4.6 instead.
The following two lemmas show that given a subset $V$ of $\mathbb{F}_{q}^{m}$, we can efficiently compute its extension degree, find a polynomial extending a function $f:V\to\mathbb{F}_{q}$, and compute a minimal Macaulay basis.
@@
For every prime power $q$ and $m\in\mathbb{N}$, there exists:
@@
Lemma 4.9 (Computing a smallest Macaulay basis).
For every prime power $q$ and $m\in\mathbb{N}$, there exists an algorithm which takes a set of $n$ points $V$ in $\mathbb{F}_{q}^{m}$ as input and gives a minimal (w.r.t. size) Macaulay basis of $\mathbb{I}(V)$ as output in the monomial basis in time $\operatorname{poly}\left(n,\log q,\binom{m+d+1}{m}\right)$, where $d$ is the extension degree of $V$.
@@
We will then construct the subsets of the minimal Macaulay basis $\mathfrak{G}_{i}$ inductively. We first define $\mathfrak{G}_{0}=\emptyset$, and then set $\mathfrak{G}_{i}$ to be any basis of the quotient space $\mathbb{I}(V)_{\leq i}/L_{i}$, where
@@
We will now show any set $\mathfrak{G}$ is a minimal Macaulay basis if and only if $\mathfrak{G}_{i}$ is a basis for the quotient space $\mathbb{I}(V)_{\leq i}/L_{i}$. To see this, if every $\mathfrak{G}_{i}$ is a spanning set for $\mathbb{I}(V)_{\leq i}/L_{i}$, then any polynomial $P\in\mathbb{I}(V)_{\leq i}$ can be written as
@@
where $c_{g}$ are constants and $\deg(h_{g}\cdot g)\leq i$. This is the condition for $\mathfrak{G}$ being a Macaulay basis, so $\mathfrak{G}$ is a Macaulay basis if and only if each $\mathfrak{G}_{i}$ spans $\mathbb{I}(V)_{\leq i}/L_{i}$.
@@
From the above, we see that a minimal Macaulay basis does not contain any degree $i$ polynomials if